# Non-linear degenerate parabolic flow equations and a finer differential structure on Wasserstein spaces

### v1

@@ 3.6.2 First Main Result: Generalised Central Limit Theorem—weak version
First, assume that $(M,g)$ is of bounded geometry. Then, Corollary 3.33 implies that the $\bigl{(}\mu^{N}\bigr{)}_{N\in\mathbb{N}}$ are uniformly $\frac{1}{p}$–Hölder continuous in the pp–Wasserstein distance. Hence, in particular, Lemma 2.4 yields that the $\bigl{(}\mu^{N}\bigr{)}_{N\in\mathbb{N}}$ are uniformlly $\frac{1}{p}$–Hölder continuous in the qq–Wasserstein distance for $1\leq q<p$. Furthermore, Lemma 2.11 implies that the $\bigl{(}\mu^{N}\bigr{)}_{N\in\mathbb{N}}$ are tight in the qq–Wasserstein distance. Thus, by Corollary 2.13, the $\bigl{(}\mu^{N}\bigr{)}_{N\in\mathbb{N}}$ are relatively compact in $\mathcal{W}_{q}(M)$. Therefore, the Theorem of Arzelà–Ascoli implies the existence of a subsequence $\mu^{M_{k}}$ which converges uniformly to a Hölder–continuous curve $\mu(t)$ with exponent $\frac{1}{p}$ in the qq–Wasserstein distance. In particular, by successive extractions, we obtain a subsequence $\bigl{(}\mu^{N_{k}}\bigr{)}_{k\in\mathbb{N}}$ satisfying $\mathcal{W}_{p-\frac{1}{k}}(\mu^{N_{k}},\mu)_{\infty}\leq\frac{1}{k}$.
Now, again by Corollary 3.33, the curves $t\mapsto\mathcal{M}_{p}(\mu^{M_{k}})$ are uniformly Lipschitz. In particular, they are uniformly bounded. Thus, the Theorem of Arzelà–Ascoli implies that there exists a subsequence $\mu^{N_{l}}\coloneqq\mu^{M_{k_{l}}}$, such that $\biggl{(}t\mapsto\mathcal{M}_{p}\bigl{(}\mu^{N_{l}}_{t}\bigr{)}\biggr{)}$ uniformly converges to a Lipschitz curve $t\mapsto M_{p}(t)$. Now, for all $k\in\mathbb{N}$,
@@
Let $\varepsilon>0$. Lemma 3.34 yields that there exists $L\in\mathbb{N}$, such that for all $l\geq L$,
$\biggl{\|}\bigl{(}t\mapsto|\mathcal{M}_{p}(\mu^{N_{l}}(t))-\mathcal{M}_{p-\frac{1}{l}}(\mu^{N_{l}}(t))|\bigr{)}\biggr{\|}_{\infty}\leq\varepsilon.$
W.l.o.g. $\frac{1}{L}\leq\varepsilon$ and for all $l\geq L$,
$\biggl{\|}\bigl{(}t\mapsto|\mathcal{M}_{p}(\mu^{N_{l}}(t))-M_{p}(t)|\bigr{)}\biggr{\|}_{\infty}\leq\varepsilon.$
Therefore, it follows for all $t$ and all $l\geq L$ that,
$\mathcal{M}_{p}(\mu^{N_{l}}(t))\geq M_{p}(t)-\varepsilon\geq\mathcal{M}_{p}(\mu(t))-\varepsilon\geq\mathcal{M}_{p-\frac{1}{l}}(\mu(t))-\varepsilon$ (3.110)
$\geq\mathcal{M}_{p-\frac{1}{l}}(\mu^{N_{l}}(t))-2\varepsilon\geq\mathcal{M}_{p}(\mu^{N_{l}}(t))-3\varepsilon.$
Hence, for all $t\in[0,T]$, $\lim_{l\to\infty}\mathcal{M}_{p}(\mu^{N_{l}}(t))=\mathcal{M}_{p}(\mu(t))$ and, thus, $\mu^{N_{l}}(t)$ converges to $\mu(t)$ with respect to the pp–Wasserstein distance by Theorem 2.1. In particular, $\bigl{(}\mu^{N_{l}}(t)\bigr{)}$ is realtively compact in $\mathcal{W}_{p}(M)$. Using Arzelà–Ascoli a last time implies, therefore, that there exists a subsequence $\bigl{(}\mu^{{\mathcal{N}}_{l}}\bigr{)}$ of $\bigl{(}\mu^{N_{l}}\bigr{)}$, which uniformly converges to $\mu$ in the pp–Wasserstein distance and $t\mapsto\mu(t)$ is Hölder–continuous with exponent $\frac{1}{p}$ in the pp–Wasserstein distance. Hence, $V[\mu^{{\mathcal{N}}_{l}}(t)]$ converges to $V[\mu(t)]$ with respect to $\mathcal{W}_{p}$ and $\overline{V}[\mu^{{\mathcal{N}}_{l}}(t)]$ converges to $\overline{V}[\mu(t)]$ with respect to $\|\leavevmode\nobreak\ \|_{C^{1}}$. Therefore, we conclude that $\mu$ solves (3.14) with initial value $\mu_{0}$.
@@
$-\int_{s}^{t}\int_{M}\int_{\mathcal{X}(M)}\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(u)]}\varphi(x)\operatorname{d}V[\mu^{N}(u)](X)\operatorname{d}\mu^{N}(u)(x)\operatorname{d}u$
@@
$\qquad\leq K(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})|t-s|\Omega(\frac{1}{N}).$
@@
$\leq\biggl{|}\int_{s}^{t}\int_{M}\int_{\mathcal{X}(M)}\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(u)]}\varphi(x)\operatorname{d}V[\mu^{N}(u)](X)\operatorname{d}\mu^{N}(u)(x)\operatorname{d}u\biggr{|}$
$+\biggl{|}\int_{s}^{t}\int_{M}\mathcal{L}_{\overline{V}[\mu^{N}(u)]}(\varphi)\operatorname{d}\mu^{N}(u)\operatorname{d}u\biggr{|}$
$+K(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})|t-s|\Omega(\frac{1}{N}).$
@@
$\biggl{|}\int_{M}\int_{\mathcal{X}(M)}\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(u)]}\varphi(x)\operatorname{d}V[\mu^{N}(u)](X)$ $\operatorname{d}\mu^{N}(u)(x)\biggr{|}$
@@
$\biggl{|}\int_{M}\varphi\operatorname{d}(\mu^{N}(t)-\mu^{N}(s))\biggr{|}\leq C\biggl{(}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})|t-s|\Omega(\frac{1}{N})$ (3.113)
$\qquad+(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty})|t-s|\biggr{)},$
@@
for some $K_{3}>0$ wthat does not depend on $N$ or $R(t)$.
@@
$\int_{M\setminus B(x_{0},R)}d(x,x_{0})\operatorname{d}\mu^{N}(t)\leq\varepsilon\qquad\text{and}\qquad\int_{M\setminus B(x_{0},R)}d(x,x_{0})\operatorname{d}\mu^{N}(0)\leq\varepsilon.$
@@
$+\int_{s}^{t}K_{2}|\nabla\varphi(x)|\operatorname{d}\mu^{N}(u)(x)\,du+|t-s|K\|\nabla^{3}\varphi\|_{\infty}\omega(\delta^{t}_{N}).$
@@
$\int_{M}K_{1}(|\nabla\varphi(x)|+|\nabla^{2}\varphi(x)|)\operatorname{d}\mu^{N}(u)(x)\leq C\mathcal{M}_{k}(\mu^{N}(s)),$ (3.117)
@@
Otherwise, there exists $k\in\mathbb{N}$ with k>p1k>p-1, for which the statement is already proved. But since the $(\mu^{N}(t))_{N,t}$ have uniformly bounded kk–th moments, they have in particular (p1)(p-1)–th moments bounded by some $K_{p-1}>0$. Now, since $\mu_{0}\in\mathcal{W}_{p}(M)$ and the $(\mu^{N}(t))_{N,t}$ are all in $\mathcal{W}_{p}(M)$ by Lemma 3.30, we proceed completely analogously as in the induction step above to prove that there exist $K_{p}>0$ such that $\mathcal{M}_{p}(\mu^{N}(t))\leq K_{p}$ for all $N,t$.
@@
Finally, we conclude that $t\mapsto\mu(t)$ is continuous in $\mathcal{W}_{p}(M)$ and that it satisfies (3.14).
@@
Take $q<p$ such that $V$ is continuous for the $\mathcal{W}_{q}$–distance. Then, Proposition 2.4 yields that $\mu(t)$ is continuous for the $\mathcal{W}_{q}$–topology on $\mathcal{W}_{p}(M)$ and for all $t\in[0,T]$, $\mu^{N_{k}}(t)$ converges to $\mu(t)$ in the $\mathcal{W}_{q}$–distance. Hence, $V[\mu^{N}(t)]$ converges to $V[\mu(t)]$ with respect to $\mathcal{W}_{p}$ and $\overline{V}[\mu^{N_{k}}(t)]$ converges to $\overline{V}[\mu(t)]$ with respect to $\|\leavevmode\nobreak\ \|_{C^{1}}$. Therefore, we conclude that $\mu$ solves (3.14) with initial value $\mu_{0}$.
@@
Finally, let us quickly justify why we call Theorem 3.35 a generalisation of the Central Limit Theorem.
Theorem 3.38 (Central Limit Theorem).
Proof.
$i:$ $\mathbb{R}^{n}\to\mathcal{X}^{2,\infty}(M)$ (3.120)
$v\mapsto X_{v}$
is an isometry with respect to the $W^{2,\infty}$–norm. Hence it is in particular measurable and $i_{*}\mu\in\mathcal{W}_{p}(\mathcal{X}^{2,\infty}(M))$. Now, define the constant vector–field probability $V:\nu\mapsto i_{*}\mu$. Note that $\overline{V}$ is constant equal to $x\mapsto\mathbb{E}(X_{1})$. Furthermore, for all $x\in\mathbb{R}^{n}$ and all $t\in[0,T]$,
$\rho^{X_{v}}_{t}(x)=x+tv.$ (3.121)
Since the $X_{i}$ are independent by hypothesis, we readily show that for all $n\in\mathbb{N}$, the distribution of $\frac{1}{\sqrt{n}}\sum_{i=0}^{n-1}(X_{i}-\mathbb{E}(X_{i}))$ is just $\bigl{(}f^{V}_{\frac{1}{n}}\bigr{)}^{\circ n}(\delta_{0})=\mu^{n}(1)$, where the $\mu^{n}$ are the Average Flow Approximation Series of Definition 3.27. Now, Theorem 3.35 gives that any subsequence of $\mu^{N}$ has a subsequence that converges to a solution of (3.14) with initial condition $\delta_{0}$ with respect to the $p^{\prime}$–Wasserstein distance.
Thus,
$\int_{\mathcal{X}(M)}$ $\mathcal{L}^{2}_{X_{v}-\mathbb{E}(X_{1})}\varphi(x)\operatorname{d}i_{*}\mu(X_{v})$ (3.123)
$=\biggl{(}\int_{\mathcal{X}(M)}(X_{v}-\mathbb{E}(X_{1}))_{i}(x)(X_{v}-\mathbb{E}(X_{1}))_{j}(x)\operatorname{d}i_{*}\mu(X_{v})\biggr{)}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}\varphi(x)$
$=\biggl{(}\int_{\mathcal{X}(M)}(v_{i}-\mathbb{E}(X_{1}))(v_{j}-\mathbb{E}(X_{1}))\operatorname{d}\mu(v)\biggr{)}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}\varphi(x)$
$=(\Sigma^{2})_{i,j}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}\varphi(x).$
Hence $\mu(t)$ solves the equation
for all $s\in[0,T]$ and all $\varphi\in\mathcal{C}^{\infty}_{c}(M)$. We know that the Gaussian diffusion $\mu(t)=\mathcal{N}(0,t\Sigma^{2})$ is a solution of (3.124). Furthermore, distributional solutions of the semi-parabolic equation (3.124) with initial condition $\mu_{0}=\delta_{0}$ are unique on $\mathbb{R}^{n}$ if we do only consider solutions that decline fast enough towards infinity. One can show that the assumption $\mu(t)$ has finite pp–th moment, for $p\geq 3$, is strong enough to get the uniqueness.
Therefore, any subsequence of $\mu^{N}$ has a subsequence that converges to $\mu(t)=\mathcal{N}(0,t\Sigma^{2})$ with respect to the $\mathcal{W}_{p^{\prime}}$–distance. Hence, $\mu^{N}$ converges to $\mu(t)=\mathcal{N}(0,t\Sigma^{2})$ with respect to the $\mathcal{W}_{p^{\prime}}$–distance. In particular, $\lim_{n\to\infty}\mu^{n}(1)=\mathcal{N}(0,\Sigma^{2})$ with respect to the $\mathcal{W}_{p^{\prime}}(\mathbb{R}^{n})$–distance. $\square$
3.7 Uniqueness results, Grönwall type estimates and Strong Generalized Central Limit Theorem
In this last section, we shall see the the solutions of (3.14) are in general not unique. However, we will prove smooth solutions, as introduced in Definition 3.28, are unique under suitable assumptions.

### v2

@@ 3.6.5 First Main Result: Generalised Central Limit Theorem—weak version
We make the stronger assumptions of the first point.
Corollary 3.36 implies that the $\bigl(\mu^{N}\bigr)_{N\in\mathbb{N}}$ are uniformly $\frac{1}{p}$–Hölder continuous in the pp–Wasserstein distance. Hence, in particular, Lemma 2.11 yields that the $\bigl(\mu^{N}\bigr)_{N\in\mathbb{N}}$ are uniformly $\frac{1}{p}$–Hölder continuous in the qq–Wasserstein distance for $1\leq q<p$. Furthermore, Lemma 2.18 implies that the $\bigl(\mu^{N}\bigr)_{N\in\mathbb{N}}$ are tight in the qq–th moment. Thus, by Corollary 2.20, the $\bigl(\mu^{N}\bigr)_{N\in\mathbb{N}}$ are relatively compact in $\mathcal{W}_{q}(M)$. Therefore, the Theorem of Arzelà–Ascoli implies the existence of a subsequence $\mu^{M_{k}}$ which converges uniformly to a Hölder–continuous curve $\mu(t)$ with exponent $\frac{1}{p}$ in the qq–Wasserstein distance. In particular, by successive extractions, we obtain a subsequence $\bigl(\mu^{N_{k}}\bigr)_{k\in\mathbb{N}}$ satisfying $\mathcal{W}_{p-\frac{1}{k}}(\mu^{N_{k}},\mu)_{\infty}\leq\frac{1}{k}$.
By Corollary 3.37, the curves $t\mapsto\mathcal{M}_{p}(\mu^{M_{k}})$ are uniformly Lipschitz. In particular, they are uniformly bounded. Thus, the Theorem of Arzelà–Ascoli implies that there exists a subsequence $\mu^{N_{l}}\coloneqq\mu^{M_{k_{l}}}$ and a Lipschitz curve $t\mapsto M_{p}(t)$, such that $\biggl(t\mapsto\mathcal{M}_{p}\bigl(\mu^{N_{l}}_{t}\bigr)\biggr)$ uniformly converges to $t\mapsto M_{p}(t)$. Now, for all $k\in\mathbb{N}$,
@@
Let $\varepsilon>0$ and $B>0$ such that for all $t\in[0,T]$, $M_{p}(t)\leq B$. Lemma 3.38 yields that there exists $L\in\mathbb{N}$, such that for all $l_{1},l\geq L$,
$\biggl\|\biggl(t\mapsto\bigl|\mathcal{M}_{p}(\mu^{N_{l}}(t))-\mathcal{M}_{p-\frac{1}{l_{1}}}(\mu^{N_{l}}(t))\bigr|\biggr)\biggr\|_{\infty}\leq\varepsilon,$
and $\bigl|B^{1-\frac{1}{Lp}}-B\bigr|\leq\varepsilon$. W.l.o.g. $\frac{1}{L}\leq\varepsilon$ and for all $l\geq L$,
$\biggl\|\biggl(t\mapsto\bigl|\mathcal{M}_{p}(\mu^{N_{l}}(t))-M_{p}(t)\bigr|\biggr)\biggr\|_{\infty}\leq\varepsilon.$
Let $t\in[0,T]$. There exists $L_{1}$ such that for all $l\geq L_{1}$,
$\biggl|\mathcal{M}_{p-\frac{1}{L}}(\mu(t))-\mathcal{M}_{p-\frac{1}{L}}(\mu^{N_{l}}(t))\biggr|\leq\varepsilon.$ (3.116)
Then, it follows for all $l\geq\max(L,L_{1})$ that
$\mathcal{M}_{p}(\mu^{N_{l}}(t))$ $\geq M_{p}(t)-\varepsilon\geq\mathcal{M}_{p}(\mu(t))-\varepsilon$ (3.117)
$\geq\mathcal{M}_{p-\frac{1}{L}}(\mu(t))^{\frac{p}{p-\frac{1}{L}}}-\varepsilon$
$\geq\mathcal{M}_{p-\frac{1}{L}}(\mu(t))-2\varepsilon\geq\mathcal{M}_{p-\frac{1}{L}}(\mu^{N_{l}}(t))-3\varepsilon$
$\geq\mathcal{M}_{p}(\mu^{N_{l}}(t))-4\varepsilon.$
Hence, for all $t\in[0,T]$, $\lim_{l\to\infty}\mathcal{M}_{p}(\mu^{N_{l}}(t))=\mathcal{M}_{p}(\mu(t))$ and, thus, $\mu^{N_{l}}(t)$ converges to $\mu(t)$ with respect to the pp–Wasserstein distance by Theorem 2.7. In particular, $\bigl(\mu^{N_{l}}(t)\bigr)$ is relatively compact in $\mathcal{W}_{p}(M)$. Using Arzelà–Ascoli one more time implies, therefore, that there exists a subsequence $\bigl(\mu^{{\mathcal{N}}_{l}}\bigr)$ of $\bigl(\mu^{N_{l}}\bigr)$, which uniformly converges to $\mu$ in the pp–Wasserstein distance and $t\mapsto\mu(t)$ is Hölder–continuous with exponent $\frac{1}{p}$ in the pp–Wasserstein distance. Hence, $V[\mu^{{\mathcal{N}}_{l}}(t)]$ converges to $V[\mu(t)]$ with respect to $\mathcal{W}_{p}$ and $\overline{V}[\mu^{{\mathcal{N}}_{l}}(t)]$ converges to $\overline{V}[\mu(t)]$ with respect to $\|\penalty 10000\ \|_{C^{1}}$. Therefore, we conclude that $\mu$ solves (MDEd) with initial value $\mu_{0}$.
@@
$-\int_{s}^{t}\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(u)]}\varphi(x)\operatorname{d}V[\mu^{N}(u)](X)\operatorname{d}\mu^{N}(u)(x)\operatorname{d}u$
@@
$\hskip 18.49988pt\leq K(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})|t-s|\Omega(\tfrac{1}{N}).$
@@
$\leq\biggl|\int_{s}^{t}\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(u)]}\varphi(x)\operatorname{d}V[\mu^{N}(u)](X)\operatorname{d}\mu^{N}(u)(x)\operatorname{d}u\biggr|$
$\hskip 18.49988pt+\biggl|\int_{s}^{t}\int_{M}\mathcal{L}_{\overline{V}[\mu^{N}(u)]}(\varphi)\operatorname{d}\mu^{N}(u)\operatorname{d}u\biggr|$
$\hskip 18.49988pt+K(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})|t-s|\Omega(\tfrac{1}{N}).$
@@
$\biggl|\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(u)]}\varphi(x)\operatorname{d}V[\mu^{N}(u)](X)$ $\operatorname{d}\mu^{N}(u)(x)\biggr|$
@@
$\biggl|\int_{M}\varphi\operatorname{d}(\mu^{N}(t)-\mu^{N}(s))\biggr|\leq C\biggl((\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})|t-s|\Omega(\tfrac{1}{N})$ (3.120)
$\hskip 18.49988pt+(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty})|t-s|\biggr),$
@@
for some $K_{3}>0$ that does not depend on $N$ or $R(t)$.
@@
$\int_{M\setminus B(x_{0},R)}d(x,x_{0})\operatorname{d}\mu^{N}(t)\leq\varepsilon\hskip 18.49988pt\text{and}\hskip 18.49988pt\int_{M\setminus B(x_{0},R)}d(x,x_{0})\operatorname{d}\mu^{N}(0)\leq\varepsilon.$
@@
$+\int_{s}^{t}K_{2}|\nabla\varphi(x)|\operatorname{d}\mu^{N}(u)(x)\,du$
$+|t-s|K\|\nabla^{3}\varphi\|_{\infty}\omega(\delta^{t}_{N}).$
@@
$\int_{M}K_{1}(|\nabla\varphi(x)|+|\nabla^{2}\varphi(x)|)\operatorname{d}\mu^{N}(u)(x)\leq C\mathcal{M}_{k}(\mu^{N}(u)),$ (3.124)
@@
Otherwise, there exists $k\in\mathbb{N}$ with k>p1k>p-1, for which the statement is already proved. But since the $(\mu^{N}(t))_{N,t}$ have uniformly bounded kk–th moments, they have in particular (p1)(p-1)–st moments bounded by some $K_{p-1}>0$. Now, since $\mu_{0}\in\mathcal{W}_{p}(M)$ and the $(\mu^{N}(t))_{N,t}$ are all in $\mathcal{W}_{p}(M)$ by Lemma 3.32, we proceed completely analogously as in the induction step above to prove that there exist $K_{p}>0$ such that $\mathcal{M}_{p}(\mu^{N}(t))\leq K_{p}$ for all $N,t$.
@@
Finally, we conclude that $t\mapsto\mu(t)$ is continuous in $\mathcal{W}_{p}(M)$ and that it satisfies (MDEd).
@@
Take $q<p$ such that $V$ is continuous for the $\mathcal{W}_{q}$–distance. Then, Proposition 2.11 and Lemma 2.18 yield that $\mu(t)$ is continuous for the $\mathcal{W}_{q}$–topology on $\mathcal{W}_{p}(M)$ and for all $t\in[0,T]$, $\mu^{N_{k}}(t)$ converges to $\mu(t)$ in the $\mathcal{W}_{q}$–distance. Hence, $V[\mu^{N}(t)]$ converges to $V[\mu(t)]$ with respect to $\mathcal{W}_{p}$ and $\overline{V}[\mu^{N_{k}}(t)]$ converges to $\overline{V}[\mu(t)]$ with respect to $\|\penalty 10000\ \|_{C^{1}}$. Therefore, we conclude that $\mu$ solves (MDEd) with initial value $\mu_{0}$.
@@
3.7 Uniqueness results, Grönwall type estimates and the Strong Generalised Central Limit Theorem
In this last section, we shall see that the solutions of (MDEd) are in general not unique. However, we will prove smooth solutions, as introduced in Definition 3.30, are unique under suitable assumptions.